# Multi-directional investigations on quiet time suprathermal ions measured by ASPEX-STEPS on-board Aditya L1

### v1

@@ Abstract
The origin, acceleration and anisotropy of suprathermal ions in the interplanetary medium during quiet periods have remained poorly understood issues in solar wind physics. To address these aspects, we derive the spectral indices for the “quiet” time suprathermal ions based on the measurements by the four directionally separated sensors that are part of the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board Aditya L1 spacecraft. Three out of four STEPS sensors - Parker Spiral (PS), Inter-Mediate (IM), Earth Pointing (EP) - are in one plane (nearly aligned with the ecliptic plane) while the fourth sensor - North Pointing (NP) – is in a mutually orthogonal plane. The energy ranges covered by the PS, IM, EP and NP sensors are 0.36-1.32 MeV, 0.14-1.22 MeV, 0.39-1.33 MeV and 0.12-1.23 MeV respectively. The quiet intervals are identified during January - November, 2024 and the derived spectral indices (differential directional flux versus energy) are found to be in the range of 2.0 for all directions in the time scale of a few days revealing isotropic nature of their distribution. Further analysis of elemental abundance ratios (${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$) during the same quiet intervals obtained from the Ultra-Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft suggests possible contributions from the leftover ions from the previous impulsive (Solar flares) and gradual events (CMEs) in the “quiet” time suprathermal ion pool.
@@ 1 Introduction
An important feature of the solar wind is the suprathermal ion tail which is manifested in the form of ions in the energy range of $\sim 10~{}\mathrm{keV~{}n^{-1}}$ to $\sim 1~{}\mathrm{MeV~{}n^{-1}}$. These suprathermal ions are ubiquitous in the heliosphere and are continuously observed to arrive at the spacecraft location from multi-directions in the interplanetary (IP) medium. The suprathermal ions act as a seed population for solar energetic particles (SEPs) accelerated by IP shocks associated with transient events in the heliosphere such as coronal mass ejections (CMEs; J. Gosling et al. 1981; G. Mason et al. 1999; M. I. Desai et al. 2003, 2004) and corotating interaction regions (CIRs; L. Fisk & M. Lee 1980; K. Chotoo et al. 2000; H. Kucharek et al. 2003; G. M. Mason et al. 2008; R. C. Allen et al. 2019). In astrophysical plasmas, particle acceleration is primarily governed by two well-established mechanisms: first-order Fermi acceleration, also known as diffusive shock acceleration (G. Krymskii, 1977; A. Bell, 1978), and second-order Fermi acceleration, originally proposed by E. Fermi (1949). Both mechanisms require the existence of a pre-accelerated seed ion population whose energies are above the injection threshold and which can be further energized within the acceleration framework. Therefore, the importance of suprathermal ion pool is undeniable to understand the generation of SEPs.
@@
The IP medium hosts a diverse reservoir of suprathermal ions, consisting of protons as well as heavier ions (4He to Fe and beyond) (M. Desai & J. Giacalone, 2016a). Studies of the composition of these suprathermal ion populations indicate that their origins can be traced to multiple sources. These sources include solar wind (e.g., M. I. Desai et al. 2003), remnants from the past transient events such as CMEs and solar flares (G. Mason et al., 1999; D. V. Reames, 2013), and interstellar pick-up ions (R. C. Allen et al., 2019). Notably, beyond 1 AU, interstellar pick-up ions have been found to contribute into the suprathermal ion pool (L. Fisk, 1976). Their abundance and energy distribution play a crucial role in shaping the overall dynamics of suprathermal ion population and contribute to the seed population for the subsequent shock acceleration in the heliosphere (A. Tylka et al., 2001). Suprathermal ions typically exhibit power-law spectra when the differential directional flux (phase space density) is plotted as a function of energy (speed). In general, a power law spectrum and its spectral index (slope of the power spectrum) correspond to a natural stochastic process. In case of suprathermal particles also, theoretical models such as pump acceleration mechanism suggests that this power-law spectrum follows a characteristic spectral index of approximately $-1.5\ (-5)$ in the differential directional flux vs. energy (phase space density vs. speed) representation (L. Fisk & G. Gloeckler, 2006, 2008, 2012, 2014) during “quiet” times. However, several studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) show that the spectral index of suprathermal ions in the 100 keV to 1 MeV energy range during “quiet” times is not universally fixed at $-1.5$, but varies over a wide range (1.22.91.22.91.2–2.91.2 – 2.9 ). Moreover, these works also reveal that, on many occasions, different elements have significantly different spectral indices during a given interval suggesting $m/q$ dependence as far as acceleration mechanisms are concerned (e.g. B. Dalal et al. 2022). Therefore, the understanding on the source and energization process(es) of suprathermal ions in the IP medium is far from complete.
@@
While dealing with acceleration of suprathermal ions in the interplanetary (IP) medium, Parker transport equation (PTE) is solved on many occasions with the assumption that suprathermal ions are isotropic in the IP medium (e.g. L. Fisk & G. Gloeckler 2008). Since suprathermal ion fluxes obtained from the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board India’s Aditya L1 are directionally resolved, experimental verification of the isotropic assumption is possible based on evaluation of spectral indices. Further, unlike previous works (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) where datasets for longer intervals are used, we use shorter intervals of data from ASPEX-STEPS for this study to minimize influence of multiple transient processes in the suprathermal ion fluxes. After selecting shorter intervals (a few days) of quiet period fluxes in 2024, we check the variations in spectral indices with directions.
@@
It is to be noted here that although directionally resolved fluxes are obtained from ASPEX-STEPS, these fluxes are not species separated and are primarily dominated by protons and alpha particles. Therefore, in order to understand the source process(es) through which these quiet period suprathermal ions are generated, elemental abundance data for the same quiet intervals are analyzed. We use the data from Ultra Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft for this purpose.
@@
Therefore, we intend to address two problems in this work. One, whether the measured suprathermal ion fluxes are directionally isotropic at L1 and two, what are the possible sources of these ions. The manuscript is organized as follows. The datasets are discussed in section 2. In Section 3, the methodology used to select the quiet intervals is discussed. The salient results are presented in Section 4. In Section 5, we discuss the results and in Section 6, the salient inferences from this work are summarized.
@@ 2 Dataset
Aditya L1 (S. Seetha & S. Megala, 2017; D. Tripathi et al., 2022) is the first observatory class mission of India to study the sun and solar wind. Aditya L1 was launched on 2nd September 2023 and placed in the halo orbit around the first Lagrange point (L1) of the Sun-Earth system on 6th January 2024. The Aditya Solar wind Particle EXperiment (ASPEX) (P. Janardhan et al., 2017; S. Goyal et al., 2018; P. Kumar et al., 2025) is one of the three in-situ experiments that measures the solar wind and energetic ions. ASPEX consists of two subsystems and the Supra-Thermal and Energetic Particle Spectrometer (STEPS) (S. K. Goyal et al., 2025) subsystem of ASPEX measures the suprathermal and energetic ions. There are six sensors in STEPS – Sun-Radial (SR), Inter-Mediate (IM), Parker Spiral (PS), North Pointing (NP), South Pointing (SP) and Earth Pointing (EP). Among the six sensors, the SR and SP sensors are experiencing light saturation issues and data from these sensors are not used in this work. As described in S. K. Goyal et al. (2025), the NP and IM units are a single-window sensor equipped with a dead layer of $0.2\,\mu\mathrm{m}$ thickness. In contrast, the PS and EP units consist of a stack of custom-built dual-window Si-PIN detector and a combined scintillator plus silicon photomultiplier (SiPM) detector assembly. The dual-window Si-PIN detector features two active regions with different dead layer thicknesses: $0.1\,\mu\mathrm{m}$ for the circular inner detector (PS-Inn and EP-Inn, with a diameter of $7\,\mathrm{mm}$) and $0.8\,\mu\mathrm{m}$ for the annular outer detector (PS-Out and EP-Out, spanning diameters from $7\,\mathrm{mm}$ to $18\,\mathrm{mm}$). Due to the extremely thin dead layer in the PS-Inn and EP-Inn sensor, it exhibits distinct spectral signatures corresponding to multiple ion species. As the deconvolution of these spectral features in the PS-Inn and EP-Inn sensor is still ongoing, this study utilizes species-integrated energetic ion flux data from the PS-Out and EP-Out sensors. Consequently, references to the PS and EP unit in this work pertain exclusively to the PS-Out and EP-Out sensors. The fluxes measured by the STEPS sensors are converted from the spacecraft frame to the GSE coordinate system. The three (PS, IM and EP) out of four remaining sensors measure primarily in the X-Y plane of the GSE coordinate system whereas the NP sensor is directed toward the +ve Z direction. STEPS measures the flux of suprathermal/energetic ions in the energy range spanning from 20 KeV/n to 6 MeV/n. We use the linear (in a log-log curve of flux vs. energy) part of the spectra to derive the spectral slopes from the fluxes. This leads to selection of ion fluxes with lower cut-off of 0.3 MeV for PS and EP sensors and 0.12 MeV for IM and NP sensors. Further, fluxes above 1.3 MeV for PS and EP or above 1.2 MeV for IM and NP are also not used for two reasons. One, fluxes in the higher energy channels are negligible during quiet periods and two, we avoid solar energetic particles ($>1.5\,\mathrm{MeV}$) in this work. For elemental abundance of suprathermal ions, data of ${}^{3}\mathrm{He}$, ${}^{4}\mathrm{He}$, C, Fe and O from the Ultra-Low Energy Isotope Spectrometer (ULEIS; E. C. Stone et al. 1998) on board the Advanced Composition Explorer (ACE) spacecraft are used. Since the objective of this work is to investigate quiet periods, the criterion for selection of quiet periods is described in the subsequent section.
@@ 3 Methodology
This section is divided into two subsections. The first subsection deals with the methodology based on which the quiet intervals are chosen and the second subsection contains the method on how spectral slopes are calculated.
3.1 Selection of quiet intervals
@@
By quiet time, we refer to the conditions in the IP medium when there is no enhancement in the suprathermal ion fluxes beyond a threshold level due to the effects of transient events like SIR/CIRs, ICMEs etc. Quiet times are selected following a heuristic statistical methodology suggested earlier by M. Dayeh et al. (2017). In this method, as a first step, hourly averaged directional differential flux is calculated based on the measurements by AL1-ASPEX-STEPS-PS, IM, EP, and NP sensors in the energy ranges of 0.36 – 1.32 MeV, 0.13 – 1.22 MeV, 0.38 – 1.32 MeV, and 0.12 – 1.23 MeV, respectively. Subsequently, the fluxes from each direction are sorted in an ascending order. These hourly-averaged and sorted fluxes are re-binned with bin size of 24 hours and the mean value as well as variance are calculated. The mean versus variance values of sorted fluxes for each sensor of AL1-ASPEX-STEPS are plotted in Figure 1.
@@
Figure 1: Mean versus Variance plots of the sorted (in ascending order) directional differential flux for Parker Spiral (PS), Earth Pointing (EP), North Pointing (NP), and Intermediate (IM) sensors of AL1-ASPEX-STEPS during the observational period from January to November 2024. The vertical red dashed line in each panel represents the threshold value of the flux below which the fluxes are considered as the quiet time flux and beyond which the fluxes are considered to be associated with transient events. The threshold values are also mentioned. The corresponding energy range is written at the top of each panel.
@@
The vertical dashed red line in each subplot corresponds to the threshold value of the mean flux beyond which the variance in flux increases monotonically. It is apparent that the mean flux values beyond the threshold value represent enhanced fluxes during transient events (CMEs, CIRs/SIRs). Consequently, the mean flux values below the threshold value represent quiet time flux of suprathermal ions in the IP medium. The quiet intervals identified for each of the detectors for the month of March 2024 is shown in Figure 2.
@@
Figure 2: Panels a, e, c and g present the temporal variations of differential directional fluxes of suprathermal ions measured by PS, IM, EP, and NP detectors of AL1-ASPEX-STEPS, respectively. Panels b, f, d, and h present the corresponding quiet time directional differential flux of suprathermal ions in PS, IM, EP, and NP detectors, respectively. The energy channels for the panels a, b, c, d and e, f, g, h are mentioned at the left side and right side of these panels respectively.
@@
In Figure 2, the original variation (panels a, e, c, g) in fluxes is juxtaposed with the quiet time variations (panels b, f, d, h) to show the typical quiet intervals. This methodology is applied for the period from January 2024 to November 2024 except for the month of April 2024 when the satellite underwent many rotations and thruster firings. Spectral indices of suprathermal ions have been calculated for all the quiet intervals obtained using the above-mentioned methodology and this is discussed in the next subsection.

### v2

@@ Abstract
The origin, acceleration and anisotropy of suprathermal ions in the interplanetary medium during quiet time have remained poorly understood issues in solar wind physics. Using measurements (in the energy range of 0.121.330.12-1.33 MeV/n) by the four detectors that are part of the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board Aditya L1 spacecraft, we address the variations in spectral indices with directions in shorter durations during January–November 2024, which coincides with the maximum phase of Solar Cycle 25. Three out of four detectors on STEPS - Parker Spiral (PS), Inter-Mediate (IM), Earth Pointing (EP) - are in one plane while the fourth detector - North Pointing (NP) – is in a mutually orthogonal plane. The derived spectral indices are found to be $1.99\pm 0.06$ regardless of directions suggesting directionally near isotropic behavior during quiet times. It is also shown that the influence of Compton-Getting effect is negligible in this assessment of directional isotropy. This result has important ramification as directional isotropy is assumed while solving the Parker transport equation to explain the acceleration of energetic particles. Further analysis of elemental abundance ratios ($^{3}\mathrm{He}/{}^{4}\mathrm{He}$, $\mathrm{Fe}/\mathrm{O}$, and $\mathrm{C}/\mathrm{O}$) during the same quiet times obtained from the Ultra-Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft suggests possible contributions from the leftover ions from the previous solar energetic particle (SEP) events in the quiet time suprathermal ion pool.
@@ 1 Introduction
An important feature of the solar wind is the suprathermal ion tail which is manifested in the form of ions in the energy range of $\sim 10~\mathrm{keV~n^{-1}}$ to $\sim 1~\mathrm{MeV~n^{-1}}$. These suprathermal ions are ubiquitous in the heliosphere and are continuously observed to arrive at the spacecraft location from multi-directions in the interplanetary (IP) medium. The suprathermal ions act as a seed population for solar energetic particles (SEPs) accelerated by IP shocks associated with transient events in the heliosphere such as coronal mass ejections (CMEs; J. Gosling et al. 1981; G. Mason et al. 1999; M. I. Desai et al. 2003, 2004; D. Chakrabarty et al. 2025) and corotating interaction regions (CIRs; L. Fisk & M. Lee 1980; K. Chotoo et al. 2000; H. Kucharek et al. 2003; G. M. Mason et al. 2008; R. C. Allen et al. 2019). In astrophysical plasmas, particle acceleration is primarily governed by two well-established mechanisms: first-order Fermi acceleration, also known as diffusive shock acceleration (G. Krymskii, 1977; A. Bell, 1978), and second-order Fermi acceleration, originally proposed by E. Fermi (1949). Both mechanisms require the existence of a pre-accelerated ion seed populations whose energies are above the injection threshold and which can be further energized within the acceleration framework. Therefore, the importance of suprathermal ion pool is undeniable to understand the generation of SEPs.
@@
The IP medium hosts a diverse reservoir of suprathermal ions, consisting of protons as well as heavier ions (4He to Fe and beyond) (M. Desai & J. Giacalone, 2016a). Studies of the composition of these suprathermal ion populations indicate that their origins can be traced to multiple sources. These sources include solar wind (e.g., M. I. Desai et al. 2003), remnants from the past transient events such as CMEs and solar flares (G. Mason et al., 1999; D. V. Reames, 2013), and interstellar pick-up ions (R. C. Allen et al., 2019). Notably, beyond 1 AU, interstellar pick-up ions have been found to contribute into the suprathermal ion pool (L. Fisk, 1976). Their abundance and energy distribution play a crucial role in shaping the overall dynamics of suprathermal ion population and contribute to the seed population for the subsequent shock acceleration in the heliosphere (A. Tylka et al., 2001). Suprathermal ions typically exhibit power-law spectra when the differential directional flux (phase space density) is plotted as a function of energy (speed). In general, a power law spectrum and its spectral index (slope of the power spectrum) correspond to a natural stochastic process. In case of suprathermal particles also, theoretical models such as pump acceleration mechanism suggests that this power-law spectrum follows a characteristic spectral index of approximately 1.5(5)-1.5\ (-5) in the differential directional flux vs. energy (phase space density vs. speed) representation (L. Fisk & G. Gloeckler, 2006, 2008, 2012, 2014) during “quiet” times. However, several studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) show that the spectral index of suprathermal ions in the 100 keV/n to 1 MeV/n energy range during “quiet” times is not universally fixed at 1.5-1.5, but varies over a wide range (1.2-1.2 to 2.9–2.9 ). Moreover, these works also reveal that, on many occasions, different elements have significantly different spectral indices during a given interval suggesting $m/q$ dependence as far as acceleration mechanisms are concerned (e.g. B. Dalal et al. 2022). Therefore, the understanding on the source and energization process(es) of suprathermal ions in the IP medium is far from complete.
@@
While dealing with acceleration of suprathermal ions in the interplanetary (IP) medium, Parker transport equation (PTE) is solved on many occasions with the assumption that suprathermal ions are isotropic in the IP medium (e.g. L. Fisk & G. Gloeckler 2008). Since suprathermal ion fluxes obtained from the Supra-Thermal and Energetic Particle Spectrometer (STEPS) of Aditya Solar Wind Particle EXperiment (ASPEX) on-board India’s Aditya L1 are directionally resolved, experimental verification of the isotropic assumption is possible based on evaluation of spectral indices. Further, unlike previous works (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) where datasets for longer intervals are used, we use shorter intervals of data from ASPEX-STEPS for this study to minimize influence of multiple transient processes in the suprathermal ion fluxes. After selecting shorter intervals (a few days) of quiet period fluxes in 2024, we check the variations in spectral indices with direction.
@@
It is to be noted here that although directionally resolved fluxes are obtained from ASPEX-STEPS, these fluxes are not species separated and are primarily dominated by protons and alpha particles. Therefore, in order to understand the source process(es) through which these quiet time suprathermal ions are generated, elemental abundance data for the same quiet time is analyzed. We use the data from Ultra Low Energy Isotope Spectrometer (ULEIS) on –board the Advanced Composition Explorer (ACE) spacecraft for this purpose.
@@
Therefore, we intend to address two problems in this work. One, whether the measured suprathermal ion fluxes are directionally isotropic at L1 and two, what are the possible sources of these ions. The manuscript is organized as follows. The datasets are discussed in section 2. In Section 3, the methodology used to select the quiet time is discussed. The results are presented in Section 4. In Section 5, we discuss the results and in Section 6, the inferences from this work are summarized.
@@ 2 Dataset
Aditya L1 (S. Seetha & S. Megala, 2017; D. Tripathi et al., 2022; N. P. Parate et al., 2025) is the first observatory class mission of India to study the Sun and solar wind. Aditya L1 was launched on 2nd September 2023 and placed in the halo orbit around the first Lagrange point (L1) of the Sun-Earth system on 6th January 2024. The Aditya Solar wind Particle EXperiment (ASPEX) (P. Janardhan et al., 2017; S. Goyal et al., 2018; J. Sebastian et al., 2025a) is one of three in-situ experiments that measures the solar wind and energetic ions. ASPEX comprises two subsystems: the Solar Wind Ion Spectrometer (SWIS) (P. Kumar et al., 2025; A. Kumar et al., 2025) and the Supra-Thermal and Energetic Particle Spectrometer (STEPS) (S. K. Goyal et al., 2025; J. Sebastian et al., 2025b). STEPS is designed to measure suprathermal and energetic ions. There are six detectors in STEPS – Sun-Radial (SR), Inter-Mediate (IM), Parker Spiral (PS), North Pointing (NP), South Pointing (SP) and Earth Pointing (EP). Among the six detectors, the SR and SP detectors are experiencing light saturation issues and data from these detectors are not used in this work. As described in S. K. Goyal et al. (2025); J. Sebastian et al. (2025b), the NP and IM units are a single-window detector equipped with a dead layer of $0.2\,\mu\mathrm{m}$ thickness. In contrast, the PS and EP units consist of a stack of custom-built dual-window Si-PIN (Silicon Positive-Intrinsic-Negative) detector and a combined scintillator plus silicon photomultiplier (SiPM) detector assembly. The dual-window Si-PIN detector features two active regions with different dead layer thicknesses: $0.1\,\mu\mathrm{m}$ for the circular inner detector (PS-Inn and EP-Inn, with a diameter of $7\,\mathrm{mm}$) and $0.8\,\mu\mathrm{m}$ for the annular outer detector (PS-Out and EP-Out, spanning diameters from $7\,\mathrm{mm}$ to $18\,\mathrm{mm}$). Due to the extremely thin dead layer in the PS-Inn and EP-Inn detector, it exhibits distinct spectral signatures corresponding to detection of multiple ion species. As the deconvolution of these spectral features in the PS-Inn and EP-Inn detector is still ongoing, this study utilizes species-integrated energetic ion flux data from the PS-Out and EP-Out detectors. Consequently, references to the PS and EP unit in this work pertain exclusively to the PS-Out and EP-Out detectors. The fluxes measured by the STEPS detectors are converted from the spacecraft frame to the Geocentric Solar Ecliptic (GSE) coordinate system. The three (PS, IM and EP) out of four remaining detectors measure primarily in the X-Y plane of Geocentric Solar Ecliptic (GSE) coordinate system whereas the NP detector is directed toward the +ve Z direction. STEPS measures the flux of suprathermal/energetic ions in the energy range spanning from 20 keV/n to 6 MeV/n. We use the linear (in a log-log curve of flux vs. energy) part of the spectra to derive the spectral slopes from the fluxes. This leads to selection of ion fluxes with lower cut-off of 0.3 MeV/n for PS and EP detectors and 0.12 MeV/n for IM and NP detectors. Further, fluxes above 1.3 MeV/n for PS and EP or above 1.2 MeV/n for IM and NP are also not used. These upper cut-offs in the energies for the detectors are chosen to ensure that the present investigation pertains to quiet time. For elemental abundance of suprathermal ions, data of $^{3}\mathrm{He}$, $^{4}\mathrm{He}$, C, Fe and O from the Ultra-Low Energy Isotope Spectrometer (ULEIS; E. C. Stone et al. 1998) on board the Advanced Composition Explorer (ACE) spacecraft are used. Since the objective of this work is to investigate quiet time, the criterion for selection of quiet time is described in the subsequent section.
@@ 3 Methodology
This section is divided into two subsections. The first subsection deals with the methodology based on which the quiet time are chosen and the second subsection contains the method on how spectral slopes are calculated.
3.1 Selection of quiet time
@@
By quiet time, we refer to the conditions in the IP medium when there is no enhancement in the suprathermal ion fluxes beyond a threshold level due to the effects of transient events like SIR/CIRs, ICMEs etc. Quiet times are selected following a heuristic statistical methodology suggested earlier by M. Dayeh et al. (2017). In this method, as a first step, hourly averaged directional differential flux is calculated based on the measurements by AL1-ASPEX-STEPS-PS, IM, EP, and NP detectors in the energy ranges of 0.361.320.36–1.32 MeV/n, 0.131.220.13–1.22 MeV/n, 0.381.320.38–1.32 MeV/n, and 0.121.230.12–1.23 MeV/n, respectively. Subsequently, the fluxes from each direction are sorted in an ascending order. These hourly-averaged and sorted fluxes are re-binned with bin size of 24 hours and the mean value as well as variance are calculated. The mean versus variance values of sorted fluxes for each detector of AL1-ASPEX-STEPS are plotted in Figure 1. For this analysis, data from January 2024 to November 2024 were taken from AL1-ASPEX-STEPS to identify the quiet periods based on the mean–variance technique.
@@
Figure 1: Mean versus Variance plots of the sorted (in ascending order) directional differential flux for Parker Spiral (PS), Earth Pointing (EP), North Pointing (NP), and Intermediate (IM) detectors of AL1-ASPEX-STEPS during the observational period from January to November 2024. The vertical red dashed line in each panel represents the threshold value of the flux below which the fluxes are considered as the quiet time flux and beyond which the fluxes are considered to be associated with transient events. The threshold values are also mentioned. The corresponding energy range is written at the top of each panel.
@@
The vertical dashed red line in each subplot corresponds to the threshold value of the mean flux beyond which the variance in flux increases monotonically. During transient events (CMEs, CIRs/SIRs), particle fluxes exhibit significant enhancements, leading to higher mean values. At the same time, these events are characterized by large fluctuations in fluxes, which results in a monotonically increasing variance with respect to mean values. In contrast, during quiet solar wind periods, the mean flux remains nearly constant with minimal variations, and consequently, no systematic increase in variance with respect to the mean is observed. We have verified that many of these transient enhancements are associated with gradual and impulsive SEP (GSEP/ISEP) events, with flux enhancements ranging from $10^{2}$ to $10^{5}$ for the PS and EP detectors, and from $10^{3}$ to $10^{6}$ for the IM and NP detectors. The characterization of some of these events is complex and will require detailed analyses, which are beyond the scope of the present work; such events will be investigated in future studies. Hence, it is apparent that the mean flux values beyond the threshold value represent enhanced fluxes during transient events (CMEs, CIRs/SIRs), whereas the mean flux values below the threshold value correspond to the quiet-time flux of suprathermal ions in the interplanetary medium. The quiet time identified for each of the detectors for the month of March 2024 is shown in Figure 2.
@@
Figure 2: Panels a, b, c and d present the temporal variations of differential directional fluxes of suprathermal ions measured by PS, EP, IM and NP detectors of AL1-ASPEX-STEPS, respectively. Panels e, f, g, and h present the corresponding quiet time directional differential flux of suprathermal ions in PS, EP, IM and NP detectors, respectively. The energy channels for the panels are mentioned at the right side of the figure.
@@
In Figure 2, the time-flux profile for March 2024 (panels a, b, c, d) is juxtaposed with the quiet time fluxes identified using the mean vs. variance technique described earlier (panels e, f, g, h). This methodology is applied for the period from January 2024 to November 2024 except for the month of April 2024 when the satellite underwent many rotations and thruster firings. Spectral indices of suprathermal ions have been calculated for all the quiet time obtained using the above-mentioned methodology and this is discussed in the next subsection.